OpenStudy (anonymous):

what is the radius? if a right cylinder has a surface area of 252pi feet and a height of 11feet?

OpenStudy (anonymous):

Right Cylinder SA = SA (top)+SA(bottom) +SA(rest of it) SA = \[SA = \pi*r^2 + \pi*r^2 + 2*\pi*rh = 252\] Solve for r : \[252 = 2*\pi*r^2 + 2*\pi*r*h\] \[2*\pi*r^2 +2*\pi*r*h -252 = 0\] Now use the quadratic formula to find r :D

OpenStudy (anonymous):

thanks:)